The return of stock A has expected value 10% and a standard deviation of 31%. The return of stock B has expected value 8% and a standard deviation of 20%. The risk-free rate is 2%. Using the Sharpe Ratio as a criterion, which stock is better?

both same?

a or b better?

impossible to tell?
Answer #1

stock A has expected return E(A) = 10%

standard deviation SD(A) = 31%

stock B has expected return E(B) = 8%

standard deviation SD(B) = 20%

Risk free rate Rf = 2%

We know that sharpe ratio = (E(r) - Rf)/SD(r)

So, sharpe ratio of A = (10 - 2)/31 = 0.2581

Sharpe ratio of B = (8 - 2)/20 = 0.3

Since sharpe ratio of B is higher, Stock B is better.